# CRS-Diff: Controllable Generative Remote Sensing Foundation Model

### v2

@@ 3.2 Image Decoupling
HED: A pre-trained deep neural network[42] is employed to predict edges and object boundaries directly from the original image, thereby capturing high-level object boundary information and low-level details.
MLSD: We use a pre-trained transformer-based model[44] to detect straight line segments in remote sensing images.
@@
Content: For content, the given image is considered as control information. Utilizing the Image Encoder in the pretrained CLIP ViT-L/14@336px[23] model, the image is transformed into feature encoding, obtaining a global embedding. This approach offers more relevant guiding conditions than text descriptions alone.
@@
Metadata: In the processing of remote sensing images, the inclusion of additional metadata, including temporal (year, month, day) and spatial (Ground Sample Distance, latitude, longitude, cloud cover) information, is considered. This information is first quantified and categorized, serving as input for the traditional diffusion model’s category guidance[29], through the labeling of these categories. Furthermore, these categories are transformed into textual descriptions with semantic information, acting as a weak textual control condition.
3.3 Multi-conditional control
@@
Additional conditional control modules are added based on the backbone structure outlined in the previous section to reconstruct the original image from optional decoupled image representation conditions. The known conditions are categorized into two types: global control and local control, with the corresponding conditional network constructed accordingly. These are then integrated with ControlNet’s control strategy[47] to enable composite control over arbitrary conditions.
Figure 2: The detailed depiction of CRS-Diff, where the Local Control structure (pink area) extracts multiple conditional image information via a feature extraction network and channels it into a trainable SD Encoder Block, subsequently integrating all information within another, frozen SD Encoder Block (blue area).
Global Control: We aim to establish a joint conditioning mechanism encompassing Caption, Content, and Metadatafig. 2. The Caption is maintained as the primary bootstrap condition, obtaining word embedding representations via a specialized CLIP text encoder, denoted as $y_{\text{c}}$. Additionally, the processed metadata and content are converted into embedding vectors through the CLIP text encoder and image encoder, respectively denoted as $y_{\text{m}}$ and $y_{\text{i}}$. Subsequently, these are transformed into eight additional feature tokens and concatenated with the caption encoded tokens via a Feed Forward Network (FFN), represented as:
$y_{\text{input}}^{\prime}=\text{Concat}\left(y_{\text{c}},w_{m}\text{FFN}_{m}(y_{\text{m}}),w_{i}\text{FFN}_{i}(y_{\text{i}})\right)$ (2)
@@
where $w_{m}$ and $w_{i}$ are the weights applied to the outputs of the FFN for $y_{\text{m}}$ and $y_{\text{i}}$, respectively. This network then integrates these encodings with the caption’s word embedding, replacing the original text encoding as the Key and Value in the cross-attention layer.
@@
Local Control: All decoupled image conditions, including sketches, segmentation masks, depth maps, road maps, and MLSD, are utilized as local control information. As shown in fig. 2, this part of CRS-Diff is based on the ControlNet, with the SD weights initially fixed to replicate the encoders and the structure and weights of the intermediate blocks. Notably, the Feature Extractor, comprising stacked convolutional neural networks, is designed to process the multiple local control information to match the dimensions of the noisy image. This processed information is then passed through a zero convolution layer, summed with the input noise, and forwarded to the replicated encoder, facilitating control over the image conditions.
@@
Given a set of image conditions $\{c_{i}\}_{i=1}^{n}$, the local control information processing can be represented as a function $\mathcal{F}$, which maps the set of conditions to a feature space that aligns with the input noise dimensions. This mapping is formally defined as :$\mathcal{F}:\{c_{i}\}\rightarrow\mathbb{R}^{d\times d}$, where $d$ represents the dimensionality of the input noise latent. The transformation leverages a series of convolutional and pooling layers to effectively capture the spatial hierarchies and semantic features of the control information, ensuring that the generated features are representative of the underlying conditions.
We integrate information from various scales through self-attention, where the transformed control information is channelled through the FDN, merged with the input noise. Here, we harmonize denoised images and local conditions via Feature Denormalization (FDN)[21] to accommodate replicated encoders of varying resolutions[50]. The merged latent space variables are subsequently introduced into a frozen encoder block (or intermediate block).By integrating local control dynamics into the generation process, we ensure the model generates images that are not only diverse but also adhere to the specified conditions with high fidelity.
3.4 Training Strategy
@@
Initial training is conducted using SD1.5 weights on a recognized text-to-image remote sensing dataset, aiming to develop a high-precision diffusion model for text-to-image generation that serves as the backbone of the ControlNet structure. This foundation enables the model to accurately guide the denoising process, leveraging a combination of multiple conditions through joint training. For both local and global control conditions, individual conditions are omitted with an independent probability of 0.5, and all conditions with a joint probability of 0.1, in accordance with Classifier-free Guidance. The dropout probability of certain conditions will be adjusted during experiments based on performance. Each metadata element is treated as a distinct bootstrap condition, with single or multiple conditions being omitted at certain probabilities, enabling the model to learn a broader array of condition combinations.
4 Experiments
4.1 Datasets
During the training of the model, we used multiple datasets:
@@ 4.1 Datasets
RSICD dataset[18]. Designed specifically for image captioning in remote sensing imagery, the RSICD contains 10,921 aerial remote sensing images accompanied by captions in natural language, sized at $224\times 224$.
@@
fMoW dataset.[1] A large-scale remote sensing image dataset, the Functional Map of the World (fMoW), includes spatio-temporal and category information for each image. Notably, only the RGB images, sized at $224\times 224$, are used, with relevant metadata extracted from a total of 110,000 images for training the multi-conditional control model.
@@
Million-AID dataset[17]. A benchmark dataset for remote sensing scene classification, Million-AID contains millions of instances, featuring 51 scene categories with 2,000 to 45,000 images per category.
4.2 Implementation details
@@
In the initial backbone model training phase, the following steps are taken: The RSICD dataset is fine-tuned over 10 epochs using a UNet network and the AdamW optimizer with a learning rate of $1\times 10^{-5}$. The input images are resized to $512\times 512$, with the model’s parameters totaling approximately 0.9 billion. For the sampling process, DDPM is utilized, setting the number of time steps to 100 and the classifier-free bootstrap scale to 7.5.
@@
During the conditional control phase: Original images from the fMoW and Million-AID datasets are organized into 200,000 text-image pairs. These images are then segmented into multiple conditional representations through annotator networks. This process involves randomly combining single or multiple conditional information, including road maps, MLSD, content, and extracted raw captions, metadata, and fine-tuning the conditional control network over 5 epochs. The AdamW optimizer, with a learning rate of $1\times 10^{-5}$, is used throughout this process, resizing both the input image and the local conditional graph to $512\times 512$. Experiments are conducted on NVIDIA GeForce RTX 4090 and NVIDIA RTX A100 GPUs, with a batch size of 8.
4.3 Evaluation metrics
We utilize four metrics to assess the effectiveness of image generation under the single text condition, namely the Inception Score (IS)[32], the Fréchet Inception Distance (FID)[8], the CLIP Score[23], and the Overall Accuracy (OA)[45] for zero-shot classification. We evaluate the performance of our proposed method and the baselines under identical settings.
  1. 1.
    Train a classifier (Resnet network[7]) utilizing the images produced by the generative model as training data.
  2. 2.
$OA=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}(y_{i}=\hat{y}_{i})$ (3)
@@ 4.3 Evaluation metrics
Figure 3: A schematic representation of remote sensing images is produced using different text-to-image generation methods based on text descriptions from the RSICD test set.
@@
Table 1: Comparisons results between our CRS-Diff and other methods in terms of four evaluation metrics (i.e., Inception score, FID score, CLIP score, and Zero-Shot classification OA on the RSICD test set). The best results are in bold
@@
Method IS $\uparrow$ FID $\downarrow$ CLIP Score ($\times$100) $\uparrow$ Zero-Shot classification OA $\uparrow$
@@
Txt2Img-MHN(VQGAN) 5.99 102.44 20.27 65.72
CRS-Diff(ours) 18.39 50.72 20.33 67.38
4.4 Comparison and Analysis
4.4.1 Single Image generation.
@@ 4.4 Comparison and Analysis
We trained CRS-Diff on the RSICD dataset using solely text as the initial input and compared it with recent SOTA methods.
@@
Qualitative analysis. In fig. 3, examples of generation using various methods, such as CRS-Diff, Txt2Img-MHN (including VQVAE and VQGAN)[45], are illustrated. Notably, our proposed CRS-Diff demonstrates the capability to generate clearer and more realistic images. For instance, when confronted with complex textual descriptions, exemplified by the phrase on the left side of the fifth line: "a square lawn and a half round lawn consist the square which is surrounded by the forest", the model proficiently deciphers the semantic content. Moreover, the model adeptly identifies the semantic information correlated with the textual descriptions, accurately reflecting this in the results. It also excels at synthesizing more appropriate images in response to shape descriptions like "square" and "long." Furthermore, CRS-Diff distinctly grasps the concept of quantity, a feature intuitively evident in its handling of numerical descriptors like "some", "many", or "two" (as seen on the left side of the first line). Additionally, our model can accurately simulate real lighting conditions, and coordinate elements such as color and texture.
@@
Quantitative analysis: We compared CRS-Diff with Attn-GAN[43], DAE-GAN[30], StrucGAN[49], DF-GAN[37], Lafite[52], DALL-E[25], Txt2Img-MHN (including VQVAE and VQGAN)[45]. The generated results were quantitatively analysed using the RSICD validation set, employing four evaluation metrics: zero-shot classification OA, initial score, CLIP score, and FID score, as delineated in table 1. Our method surpassed the baseline in three metrics and achieved third place in CLIP score. Intriguingly, the performance in CLIP score was not as high as anticipated, possibly due to the specificities of the CLIP model employed in our evaluation. Nevertheless, the proposed CRS-Diff demonstrates excellent performance in controllability and generation quality, meeting the demands of practical applications like urban planning and laying a foundation for future research in controlled generation.

### v3

@@ III-B Image Decoupling
HED (Holistically-nested Edge Detection): A pre-trained deep neural network [46] is employed to predict edges and object boundaries directly from the original image, thereby capturing high-level object boundary information and low-level details.
MLSD (Multiscale Line Segment Detection): We use a pre-trained transformer-based model [47] to detect straight line segments in remote sensing images.
@@
Content: For content, the given image is considered as control information. Utilizing the Image Encoder in the pre-trained CLIP ViT-L/14@336px [41] model, the image is transformed into feature encoding, obtaining a global embedding. This approach offers more relevant guiding conditions than text descriptions alone.
@@
Metadata: In processing the RS image, it is crucial to incorporate additional metadata, such as temporal (year, month, day) and spatial (ground sampling distance, latitude, longitude, cloud cover) information. This information is first quantified and categorized, serving as input for the traditional diffusion model’s category guidance [6], through the labeling of these categories. Additionally, these metadata are transformed into sequence tokens, which are incorporated into the text encoding as weak text control conditions.
III-C Multi-conditional fusion
@@
Based on the backbone structure outlined in the previous section, additional conditional control modules are added to reconstruct the original image from the decoupled image representation conditions. This process trains the model’s multi-conditional generation capability. The known conditions are categorized into three types: image conditions, text conditions, and metadata. For each type of condition, we explore the corresponding condition injection methods and construct a feature extraction network that meets the requirements. The obtained condition features are then integrated with the ControlNet control strategy [30] through feature fusion to achieve composite control of arbitrary conditions, as illustrated in Fig. 3 and Fig. 4.
Text Conditional Fusion: We aim to establish a joint conditional bootstrapping mechanism that includes captions, content, and metadata, as illustrated in Fig. 4. The text description, as the main bootstrap condition, is represented as $y_{\text{t}}$ by obtaining a word embedding through a specialized CLIP text encoder. Considering that text input in T2I generation is a natural method, We map the metadata to an interval with a fixed range and encode it into a corresponding number of metadata encodings through a linear layer, with the result denoted as $y_{\text{m}}$. Additionally, we introduce an image encoder to encode the content image and then use a Feed-Forward Network (FFN) to connect the feature vectors with the prompt encoding symbols, represented as:
$y^{\prime}=\text{Concat}\left(y_{\text{t}},w_{\text{c}}\cdot\text{FFN}(y_{\text{c}}),w_{\text{m}}\cdot\text{MLP}(y_{\text{m}}\right)),$ (3)
@@
where $w_{c}$ and $w_{m}$ are the weights applied to the outputs of $y_{\text{c}}$ and the $y_{\text{m}}$, respectively. This network subsequently integrates these encodings with the caption’s word embedding, replacing the original input tokens as the Key and Value in the cross-attention layer.
@@
Image Conditional Fusion: All decoupled image conditions, including Sketch, Segmentation mask, Depthmap, HED, Road map, and MLSD, are utilized as local control information. As shown in Fig. 3, this part of CRS-Diff is based on the ControlNet, with the SD weights initially fixed to replicate the encoders and the structure and weights of the intermediate blocks. It is worth noting that the feature extractor, consisting of stacked convolutional neural networks, contains multiple feature condition information in the feature map. We then fuse the latent features with the denoising latent variables through a feature fusion network. Consistent with Uni-ControlNet [39], we perform the feature injection step in four downsampled ResNet blocks within the U-Net structure of the diffusion model.
@@
Specifically, given a set of image conditions $\{c_{i}\}_{i=1}^{n}$, image conditional information processing can be represented as a function $\mathcal{F}$, which maps the set of conditions to a feature space that equations with the input noise dimensions. This mapping is formally defined as :$\mathcal{F}:\{c_{i}\}\rightarrow\mathbb{R}^{c\times d\times d}$, where $d$ represents the dimensionality of the latent feature map. The transformation leverages a series of convolutional and pooling layers to effectively capture the spatial hierarchies and semantic features of the control information, ensuring that the generated features are representative of the underlying conditions. Based on this, we resample the obtained feature map to the same dimension as the current latent variable and use Attention Feature Fusion (AFF) [52] to achieve a better fusion of the noise and feature images. This process replaces the inputs of the original residual block of the U-Net, allowing us to achieve multi-conditional information injection, represented as:
$\mathbf{z}^{\prime}=\text{AFF}(\mathbf{z},\text{Resample}(\mathcal{F}(\{c_{i}\}),\text{dim}(\mathbf{z}))),$ (4)
where $\mathbf{z}$ is the current latent variable, $\{c_{i}\}$ is the set of image conditions, ($\mathcal{F}(\{c_{i}\})$ extracts the feature map from the conditions, and $\text{Resample}(\cdot)$ adjusts the feature map dimensions to match $\mathbf{z}$, $\text{AFF}(\cdot)$ fuses the resampled feature map with $\mathbf{z}$, $\mathbf{z}^{\prime}$ is the fused latent variable. This formula replaces the inputs of the original residual block of the U-Net, allowing us to achieve multi-conditional information injection.
III-D Training Strategy
@@
Our CRS-Diff employs a two-stage training strategy to train the native Stable Diffusion (SD) architecture and the ControlNet architecture within its framework, respectively. Initial training is conducted using SD 1.5 weights on a text-to-image RS dataset, aiming to develop a high-precision diffusion model for text-to-image generation that serves as the backbone of the ControlNet structure. This foundation enables the model to accurately guide the denoising process, leveraging a combination of multiple conditions through joint training. For both text and image control conditions, individual conditions are omitted with an independent probability of 0.5, and all conditions with a joint probability of 0.1, in accordance with Classifier-free Guidance. The dropout probability of certain conditions will be adjusted during experiments based on performance. Each condition element is treated as a distinct bootstrap condition, with single or multiple conditions being omitted at certain probabilities, enabling the model to learn a broader array of condition combinations.
IV Experiments
IV-A Datasets
During the training stage, our CRS-Diff used the following datasets:
@@ IV-A Datasets
  • RSICD dataset [44]: This dataset is designed specifically for image captioning in remote sensing imagery, and it contains 10,921 aerial remote sensing images accompanied by captions in natural language, sized at $224\times 224$.
@@
  • fMoW dataset [45]: This dataset is a large-scale remote sensing image dataset, and fMoW includes the spatio-temporal and category information for each image. Notably, only the RGB images, sized at $224\times 224$, are used, with relevant metadata extracted from a total of 110,000 images for training the multi-conditional control model.
@@
  • Million-AID dataset [53]: This dataset is a benchmark dataset for remote sensing scene classification, and contains millions of instances, featuring 51 scene categories with 2,000 to 45,000 images per category.
IV-B Implementation details
@@
In the initial backbone model training phase, the following steps are taken: The RSICD dataset is fine-tuned over 10 epochs using the U-Net and the AdamW optimizer with a learning rate of $1\times 10^{-5}$. The input images are resized to $512\times 512$, with the model’s parameters totaling approximately 0.9 billion. For the sampling process, DDPM is utilized, setting the number of time steps to 100 and the classifier-free bootstrap scale to 7.5.
@@
During the conditional control phase, original images from the fMoW and Million-AID datasets are organized into 200,000 text-image pairs. These images are then segmented into multiple conditional representations through annotator networks. This process involves randomly combining single or multiple conditional information, including road maps, MLSD, content, and extracted raw captions, metadata, and fine-tuning the conditional control network over 5 epochs. The AdamW optimizer, with a learning rate of $1\times 10^{-4}$, is used throughout this process, resizing both the input image and the local conditional graph to $512\times 512$. Experiments are conducted on NVIDIA GeForce RTX 4090 and NVIDIA RTX A100 GPUs, with a batch size of 8.
@@
Figure 5: Visual comparison of different text-to-image generation methods based on text descriptions on the RSICD test set.
@@
TABLE II: Comparisons results between our CRS-Diff and other methods regarding four evaluation metrics (i.e., Inception score, FID score, CLIP score, and Zero-Shot classification OA on the RSICD test set).
@@
Method Inception Score $\uparrow$ FID Score $\downarrow$ CLIP Score ($\times$100) $\uparrow$ Zero-Shot classification OA $\uparrow$
@@
Txt2Img-MHN (VQGAN) [55] 5.99 102.44 20.27 65.72
RSDiff [17] 7.22 66.49 - -
SD 1.5 (original) [6] 6.33 167.19 19.92 32.81
SD 1.5 (fine-tune) 16.48 60.14 20.98 63.37
CRS-Diff (Ours) 18.39 50.72 20.33 69.21
IV-C Evaluation metrics
For text-to-image generation tasks, we utilize four metrics to assess the effectiveness of image generation under the single text condition, namely the Inception Score (IS) [56], the Fréchet Inception Distance (FID) [57], the CLIP Score [41], and the Overall Accuracy (OA) [55] for zero-shot classification. We evaluate the performance of our proposed method and the baselines under identical settings.
$IS=\exp\left(\mathbb{E}_{\mathbf{x}\sim p_{g}}\left[D_{\text{KL}}(p(y|\mathbf{x})||p(y))\right]\right),$ (5)
where $p(y|\mathbf{x})$ is the conditional label distribution given an image $\mathbf{x}$ and $p(y)$ is the marginal label distribution. $D_{\text{KL}}$ denotes the Kullback-Leibler divergence, and $p_{g}$ represents the distribution of generated images.
$FID=||\mu_{r}-\mu_{g}||^{2}+\text{Tr}(\Sigma_{r}+\Sigma_{g}-2(\Sigma_{r}\Sigma_{g})^{\frac{1}{2}}),$ (6)
where $\mu_{r}$ and $\Sigma_{r}$ are the mean and covariance of the real images’ features, and $\mu_{g}$ and $\Sigma_{g}$ are the mean and covariance of the generated images’ features.
  1. 1.
    Train a classifier (e.g., ResNet [58]) by utilizing the images produced by the generative model as training data.
  2. 2.
$OA=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}(y_{i}=\hat{y}_{i}),$ (7)
For conditional image generation tasks, we utilized three metrics (i.e., SSIM, mIoU, and CLIP score) to evaluate the performance of conditional image generation. We evaluated our proposed method and the baseline methods under identical settings. The calculation formulas for the SSIM and mIoU metrics are as follows:
$SSIM(x,y)=\frac{(2\mu_{x}\mu_{y}+c_{1})(2\sigma_{xy}+c_{2})}{(\mu_{x}^{2}+\mu_{y}^{2}+c_{1})(\sigma_{x}^{2}+\sigma_{y}^{2}+c_{2})},$ (8)
where $\mu_{x}$ is the mean of image $x$, $\mu_{y}$ is the mean of image $y$, $\sigma_{x}^{2}$ is the variance of image $x$, $\sigma_{y}^{2}$ is the variance of image $y$, $\sigma_{xy}$ is the covariance of images $x$ and $y$, and $c_{1}$ and $c_{2}$ are constants to stabilize the division.
$mIoU=\frac{1}{n}\sum_{i=1}^{n}\frac{|P_{i}\cap G_{i}|}{|P_{i}\cup G_{i}|},$ (9)
where $n$ is the number of classes, $P_{i}$ is the predicted region for class $i$, and $G_{i}$ is the ground truth region for class $i$.
IV-D Comparison and Analysis
IV-D1 Text-to-Image generation
@@
We trained CRS-Diff on the RSICD dataset using solely text as the initial input and compared it with recent state-of-the-art (SOTA) methods.
@@
Qualitative analysis. In Fig. 5, generated images using various methods, such as CRS-Diff, Txt2Img-MHN (including VQVAE and VQGAN) [55], are illustrated. Notably, our proposed CRS-Diff demonstrates the capability to generate clearer and more realistic images compared with other methods. For instance, when confronted with complex textual descriptions, exemplified by the phrase on the left side of the fifth line: “a square lawn and a half round lawn consist the square which is surrounded by the forest”, the model proficiently deciphers the semantic content. Moreover, the model adeptly identifies the semantic information correlated with the textual descriptions, accurately reflecting this in the results. It also excels at synthesizing more appropriate images in response to shape descriptions like “square” and “long.” Furthermore, CRS-Diff distinctly grasps the concept of quantity, a feature intuitively evident in its handling of numerical descriptors like “some”, “many”, or “two” (as seen on the left side of the first line). Besides, our model can accurately simulate real lighting conditions, and coordinate elements such as color and texture.
Figure 6: Visual comparison results of generated RS images from metadata (i.e., month and cloud cover) and image content. Except for the image content condition, metadata is used together with the textual descriptions.
@@
Quantitative analysis. We compared CRS-Diff with Attn-GAN [22], DAE-GAN [23], StrucGAN [24], DF-GAN [25], Lafite [26], DALL-E [54], Txt2Img-MHN (including VQVAE and VQGAN) [55], RSDiff [17] and SD (fine-tuned on sd1.5) [6]. The generated results are quantitatively analysed using the RSICD test set, employing four evaluation metrics: zero-shot classification OA, initial score, CLIP score, and FID score, as delineated in Tab. II. Our method surpassed the baseline in three metrics and achieved second place in the Inception Score. Intriguingly, the performance in the CLIP score is not as high as anticipated, possibly due to the specificities of the CLIP model employed in our evaluation. Nevertheless, the proposed CRS-Diff demonstrates excellent performance in controllability and generation quality, meeting the demands of practical applications like urban planning and laying a foundation for future research on controllable generation.
IV-D2 Single-condition image generation
Except for the text condition, CRS-Diff can support more conditions to guide the model towards generating more refined images.
TABLE III: Results comparison (i.e., SSIM, mIoU, and CLIP Score metrics.) between our CRS-Diff with ControlNet and Uni-ControlNet on the RSICD test set in terms of seven image-based conditions.
Method HED MLSD DepthMap Sketch Seg Road Content
(SSIM) (SSIM) (SSIM) (SSIM) (mIoU) (mIoU) (CLIP Score)
ControlNet [30] 0.4363 0.8345 0.8003 0.2895 0.3841 0.3458 0.7789
Uni-ControlNet [39] 0.4475 0.7311 0.8743 0.5288 0.3041 0.3351 0.8453
CRS-Diff (Ours) 0.4548 0.7455 0.8705 0.5676 0.2982 0.3787 0.8620
TABLE IV: Results comparison (i.e., FID metric.) between our CRS-Diff with ControlNet and Uni-ControlNet on the RSICD test set in terms of seven image-based conditions.
Method HED MLSD DepthMap Sketch Seg Road Content
ControlNet [30] 31.26 62.21 58.27 66.54 72.65 102.57 63.98
Uni-ControlNet [39] 44.76 59.89 47.10 58.73 79.15 93.08 53.97
CRS-Diff (Ours) 30.18 55.75 54.40 68.29 70.05 94.16 44.93
Qualitative analysis. Fig. 6 shows the visual comparison results of generated RS images from a single metadata condition. To mitigate the risk of inaccurate results due to conflicts between the content condition and the semantics of the textual description, we added textual guidance to all conditions except the content condition. The content condition provides richer semantic information. Metadata control proved more challenging, so we chose salient attributes such as month and cloud cover to offer more granular control.
Fig. 7 shows the visual comparison results of generated RS images from a single image condition. In HED and Sketch, intuitive image control is achieved by restricting the boundary and contour information of the generated image. Features, e.g., segmentation masks and roadmaps, can provide richer semantic information, which can be efficiently interpreted by CRS-Diff to influence the generated outputs. Conversely, CRS-Diff generates images with clear texture details and coherent scene relationships, enabling better comprehension even in areas not covered by feature conditions.